A 0.170 kg cue ball strikes a 0.160 kg 11-ball at 3.54 m/s. The cue ball stops passing all of its momentum to the 11-ball. What is the velocity of the 11-ball after impact?
1
Expert's answer
2020-10-30T13:18:20-0400

The total momentum of the cue ball (before the impact) was:


"p_c = m_cv_c = 0.170kg\times 3.54m/s = 0.6018\space kg\cdot m/s"

The cue ball passed all of its momentum. Thus, the momentum of the 11-ball (after impact) is:


"p_{11} = m_{11}v_{11} = p_c"

where "m_{11} = 0.16kg" is the mass of the 11-ball and "v_{11}" is its required velocity after impact. Expressing "v_{11}" from the last equality, obtain:


"v_{11} = \dfrac{p_c}{m_{11}} = \dfrac{0.6018}{0.16} = 3.76125 \space m/s"

Answer. 3.76125 m/s.